Kathy bicycles 5km/h faster than Javier. In the same time it takes Javier to bicycle 54 km, Kathy can bicycle 69 km. How fast does each bicyclist travel?
Javier travels _____km/h.
Kathy travels ______km/h.

Answer by htmentor(1343)

Let x = Javier's speed
Then x+5 = Kathy's speed
Since t = d/v, and the times are the same, we can write
54/x = 69/(x+5)
Solve for x:
54(x+5) = 69x
15x = 270
x = 18
So Javier's speed is 18 km/h, Kathy's speed is 23 km/h